# Artificial Intelligence Exploring the Patent Field

### v1

@@ 4.2 Curated Data Collections
TREC-CHEM [^4] TREC-CHEM: https://trec.nist.gov/data/chem-ir.html is a part of the Text REtrieval Conference (TREC) series, specialized on chemical information retrieval. It contains patent datasets that are rich in chemical information. For example, the dataset from TREC-CHEM 2009 contains 2.6 million patent files registered at European Patent Office, United States Patent and Trademark Office, and World Intellectual Property Organization [35].
5 Methods
Automation for patent tasks normally involves the two steps of feature extraction [^5] Features in the machine learning context refer to the measurable attributes of the target object, which is transformed to numerical vectors as input for model training. Patent features in the legal background introduced in Section 7.4.1 are elements that an invention has to include to become the invention. and model development. Patent documents consist of multiple aspects, such as texts, citations, and images. These data can be extracted and transformed into numerical values to represent patent documents. Then, specialized models are developed for different types of patent tasks based on these features. Notably, deep-learning models may have already included feature extraction functions, so there is usually no need for explicit feature extraction before applying deep-learning methods. In this section, we first introduce various feature representations in Section 5.1. Subsequently, we summarize widely used models in Section 5.2, including math-based functions, traditional machine learning, deep learning, and pre-trained language models. Moreover, we illustrate other useful algorithms and approaches for patent tasks in Section 5.3, such as domain adaptation and knowledge graph.
Figure 4: Overview of commonly used methods in patent-related tasks.
5.1 Feature Representations
5.1.1 Text Feature
Traditional Text Features. Bag-of-words (BoW) represents text by the frequency of each word, disregarding grammar and word order [23]. Term frequency-inverse document frequency (TF-IDF) weighs the frequency of each word by its inverse frequency in the corpus, reducing the importance of common words, such as “the” [36]. In these approaches, words are treated as discrete atomic symbols so that no relationships between words are considered. Due to the disadvantages, frequency-based methods have recently been replaced by the following embedding approaches.
Word Embedding. In natural language processing (NLP), embeddings are dense numerical vectors that represent words, phrases, or entire sentences. Word embedding transforms each word into a fixed-length numerical vector, capturing more word meanings and semantic relationships. Word2Vec [37] and global vectors for word representation (GloVe) [38] are static word embeddings, which first introduce the concept of pre-training to natural language processing tasks. The embedding is pre-trained on large unlabeled texts to capture general patterns, which map each word to a vector. Words with similar meanings would have less distance. These embeddings are easy to train and can capture the relationship between words. Nonetheless, they cannot retain word contexts or solve the problem of one word with multiple meanings. Dynamic word embeddings can solve some of these issues. For example, embeddings from language models (ELMo) [39] uses bi-directional methods [40] to capture additional context information, improving the performance of word embeddings.
Domain-specific Embedding. Word embeddings are normally trained on general corpora, such as Wikipedia. However, patent texts are highly technical and contain much jargon, leading to different text features compared to mundane texts. Additionally, patent language is highly artificial, including the terminology. This is sometimes due to intentional design, but in most cases, it is a consequence of a patent professional drafting the text who is not working in the specific field but has good general natural science and engineering knowledge. They create and delineate the requisite nomenclature, including terms that may not be found in any dictionary. This process facilitates the effective use of these terms in patent descriptions. Therefore, general text embeddings or language models can fail to capture patent characteristics if some important terms are not in the training data or entire neologisms. Thus, the model can be trained on patent texts to obtain domain-specific embeddings to improve the in-domain performance [41]. This method works for not only word embeddings but also other language models, which Section 5.3.2 illustrates.
Sentence/Paragraph Embedding. Apart from word embeddings, larger blocks of text, such as sentences, paragraphs, or entire documents can be transformed into fixed-size vectors using sentence/paragraph embedding. Doc2Vec, also known as paragraph vector, is an extension of Word2Vec, which learns fixed-length text feature representations from variable-length texts [42]. In addition, pre-trained language models (PLMs), such as bidirectional encoder representations from transformers (BERT) [43], include mechanisms for sentence/paragraph embedding. While static word embeddings, such as Word2Vec [37], provide single representations for each word regardless of its context, BERT embedding for texts changes based on the other words in the sentence, allowing the model to capture nuances in meaning that depend on context. We introduce details of pre-trained language models in Section 5.2.4. Pre-trained language models are powerful in capturing intricate contextual information, but they can be computationally intensive. Hence, researchers proposed Sentence-BERT to improve efficiency while maintaining accuracy from BERT [44].
5.1.2 Numerical Feature
Numerical features are quantitative values that are usually discrete in the patent domain. Commonly used numerical features in patents are metadata, such as the number of citations, claims, and assignees [45]. Furthermore, researchers used one-hot vectors to represent categorical features, such as classification codes [46].
5.1.3 Graph Feature
Graphs consist of nodes (or vertices) and edges (connections between nodes). Thus, graphs can represent complex connections between patents, such as citation and classification relationships. In the case of patents, nodes are attributes of patents, and edges are the relationship between patents. Graph embedding enables complex graph structures to be represented in a format suitable for analysis. For example, researchers built graph embedding for patent valuation based on patent citation networks and attributes [47], and used classification code co-occurrence information for patent retrieval [48].
5.1.4 Image Feature
Images consist of pixels, so it is simple to convert images to numerical values by using intensity values for image feature representation. Histogram of oriented gradients (HOG) and scale-invariant feature transform (SIFT) are representative of traditional image features used in patents [21; 49]. HOG describes the distribution of intensity gradients and edge directions in an image, commonly used in object detection [50]. SIFT detects and describes local features in images, which is robust to scaling and rotation [51]. Deep learning models, such as convolutional neural networks (CNNs) [52], automatically detect image features from pixel values of the full raw image, so feature extraction is not needed beforehand [53].
5.2 Models
5.2.1 Math Function
Pre-defined rules or mathematical formulas that do not require training can process patents and solve patent-related tasks with overall lower computational cost. This sub-field shares some tools with statistics and bibliometry.
Pre-defined Rules. Researchers manually pre-define a set of rules for the task to fulfill specific aims. For example, Chiarello et al. designed rule-based systems to identify and extract design affordances from patents [54]. Additionally, researchers used pre-defined indicators for novelty analysis. Plantec et al. calculated novelty scores by weighting the importance of different indicators, such as the number of citations [55]. Moreover, Ferraro et al. proposed rules for patent text simplification, which segmented complex texts into simple and easy-to-understand components [56].
Overall, this approach is easy to implement and cheap in computation. However, the performance may not be satisfactory compared to latest machine learning approaches for several reasons. The methods are based on substantial external design, intuition, and assumptions. For example, pre-defined rules may not cover all critical aspects of the task. Furthermore, the human-defined weights for each indicator may be biased.
Similarity. Similarity methods determine how similar two or more entities are. After patent documents are represented by feature vectors, similarity metrics quantify the similarity between patents. Researchers commonly use this method in patent retrieval tasks to retrieve relevant patent documents based on similarities [23]. Similarity is one of the mainstream methods for patent retrieval and demonstrating promising performance. Similarity methods can also serve for patent novelty prediction. Researchers sometimes consider a new patent more novel if it has lower similarity scores compared to existing patents [25]. Although researchers demonstrated the feasibility of using similarity for patent novelty prediction, it is less common and effective compared to other approaches introduced in Section 7.4.3. Another study proposed to use similarity for patent classification. Similar patents may have a higher probability of becoming the same category [57]. However, the performance of using similarity cannot surpass the counterparts of developing classification models directly.
In addition, the choice of similarity measures will impact the performance of algorithms. Hence, it is critical to choose appropriate similarity methods that best fit the problem. For example, Euclidean distance measures the straight distance between two points. It is effective in physical spaces but may not be suitable in high-dimensional spaces due to the curse of dimensionality. Cosine similarity measures the cosine of the angle between two vectors, which is particularly popular in text analysis for comparing the similarity of documents. Other similarity metrics include Jaccard similarity, Manhattan distance, Hamming distance, etc.
Algorithms. Various formula-based algorithms for patent tasks do not require explicit training. BM25 is a ranking function that estimates the relevance of documents to a given search query used in information retrieval [58]. The relevance score is calculated based on the frequency of query terms in each document, adjusted for the length of the document and the average document length across the whole document set. Thus, researchers used this algorithm for patent retrieval [6]. Since BM25 is based on word frequency, modern deep-learning methods that leverage semantic information for patent retrieval show better prospects.
K-nearest neighbors (KNN) is a simple and non-parametric algorithm used for classification and regression [59]. When making predictions, KNN searches through the training dataset to find the K-nearest data points (neighbors) based on a distance metric (e.g., Euclidean distance). The class of majority neighbors determines the final result. While there is no explicit training phase, the algorithm still requires the training data during the prediction phase. In the patent domain, Bekamiri et al. used KNN for patent classification, but the performance is not satisfactory compared to deep learning methods [57].
[…]

### v2

@@ 4.2 Curated Data Collections
TREC-CHEM [^11] TREC-CHEM: https://trec.nist.gov/data/chem-ir.html is a part of the Text REtrieval Conference (TREC) series, specialized on chemical information retrieval. It contains patent datasets that are rich in chemical information. For example, the dataset from TREC-CHEM 2009 contains 2.6 million patent files registered at the European Patent Office, United States Patent and Trademark Office, and World Intellectual Property Organization (Lupu et al., 2009).
5 Patent Analysis Tasks
Structure Label Description
Sub-class F02D Controlling combustion engines
Figure 5: Illustration of three evaluation methods for patent classification, where 1, 2, …, n are the top-n predictions, MC stands for main class, and IC is incidental class (Fall et al., 2003).
Patent analysis tasks focus on understanding and usage of patents. We divide patent analysis tasks into four main types, subject classification, information retrieval, quality assessment, and technology insights. Patent subject classification (Section 5.1) is one of the most widely studied topics in the patent domain, where the categories of patents are predicted based on their content. Information retrieval tasks consist of two sub-tasks, specifically patent retrieval (Section 5.2) and information extraction (Section 5.3). While patent retrieval aims at retrieving target documents from databases, information extraction focuses on extracting desired information from patent texts for further applications. Quality assessment refers to evaluating the quality of patents, which includes novelty prediction (Section 5.4), granting prediction (Section 5.5), litigation prediction (Section 5.6), and patent valuation (Section 5.7). As novelty is essential for patents, early prediction of novelty and auxiliary methods for novelty assessments could ensure patent quality before filing and improve efficiency. Granting prediction forecasts whether the patent office will grant a patent application. The process involves further aspects beyond novelty and inventiveness, such as formal requirements of the language, figures, and documents. A low-quality patent is likely to be rejected by the examiner. Litigation prediction measures the odds that the file may at some point become the subject of litigation. For example, patents with unclear or ambiguous claims or very scarce descriptions tend to be more likely to cause litigation cases. Patent valuation refers to measuring the value of the patents, which is a reflection of patent quality and scope. Technology insights are the usage of patents, consisting of technology forecasting (Section 5.8) and innovation recommendation (Section 5.9). Since patents contain extensive emerging technology information, researchers can analyze patents to predict future technological development trends or suggest new ideas for technological innovation.
5.1 Automated Subject Classification
5.1.1 Task Definition of Subject Classification
5.1.2 Classification Scheme
5.1.3 Evaluation
Most research uses one or more evaluation measures originated from Fall et al. (2003). As Fig. 5 illustrates, there are three different evaluation methods, namely top prediction, top N guesses, and all categories. Top prediction only checks whether the top-1 prediction is the same as the main class. In top N guesses, the result is successful if one of the top-n predictions matches the main class, which is more flexible compared to the top-1 prediction. On the other hand, the all-categories method checks whether the top-1 prediction is included in the set of the main class and all incidental classes.
5.1.4 Methodologies for Automated Subject Classification
We summarize the mainstream techniques for patent classification and categorize them into three types, including feature extraction and classification, fine-tuning transformer-based language models, and hybrid methods (Fig. 6).
Figure 6: Three mainstream methods for patent classification
Feature Extraction and Classifier. Researchers extract various features from patent documents and adopted a classifier for prediction based on the features (Shalaby et al., 2018; Hu et al., 2018; Abdelgawad et al., 2019; Zhu et al., 2020).
Most of the research used text content for prediction. The commonly used text representation is word embedding. Roudsari et al. (2021) compared five different text embedding approaches, including bag-of-words, GloVe, Skip-gram, FastText, and GPT-2. Bag-of-words is based on word frequency and ignores semantic information. GloVe, Skip-gram, and FastText are deep learning methods that capture word meaning but not word contexts. In contrast, GPT-2 is based on transformer architecture and is capable of capturing complex contextual information. The results demonstrated that GPT-2 performed the best with a precision of 80.52%, indicating that transformer-based embedding approaches may perform better than other traditional or deep-learning embeddings. In addition, word embeddings are typically pre-trained on heterogeneous corpora, such as Wikipedia, making them lack domain awareness. Therefore, Risch and Krestel (2019) trained a domain-specific word embedding for patents and incorporated recurrent neural networks for patent classification. This method improved the precision by 4% compared to normal word embedding, demonstrating the effectiveness of domain adaptation. Apart from word embeddings, researchers also adopted sentence embedding for feature extraction (Bekamiri et al., 2021). Other research calculated the semantic similarity between patent embeddings obtained from Sentence-BERT (Reimers and Gurevych, 2019) and used the k-nearest neighbors (KNN) method for classification with a precision of 74%. Although the performance was not strong enough, the use of similarity and KNN provided a different view of the classification approach.
Paper
Data
source
Dataset
size
Parts used
Number of
classes
Methods Metric
Results
(Top prediction)
Feature Extraction and Classification
(Risch and Krestel, 2019) USPTO-2M 1.95, 0.05 Title, abstract 637
Domain-specific
word embedding + RNN
P 53.00
(Li et al., 2018) USPTO-2M 1.95, 0.05 Title, abstract 637 Skip-gram + CNN P 73.88
(Roudsari et al., 2021) USPTO 0.24, 0.04 Title, abstract 89 GPT-2 + CNN P 80.52
(Bekamiri et al., 2021) USPTO 1.37, 0.12 Title, abstract, claims 663 SBERT + KNN P 74.00
(Jiang et al., 2021) USPTO 0.35, 0.05 Image 8 CNN A 54.32
Fine-tuning Transformer-based Language Models
(Lee and Hsiang, 2020b) USPTO-2M 1.95, 0.05 Title, abstract 637 BERT P 81.75
(Haghighian et al., 2022) USPTO-2M 1.95, 0.05 Title, abstract 544 XLNET P 82.72
(Christofidellis et al., 2023) USPTO 0.24, 0.04 Title, abstract 89
SciBERT + Domain-adaptive
training + adapters
P 84.53
(Risch et al., 2020b) USPTO-2M 1.95, 0.05 Title, abstract 632 Transformer A 56.70
Hybrid Methods
(Jiang et al., 2022a) USPTO 0.72, 0.08 Title, abstract, image 622 Multimodal A 65.50
(Zhang et al., 2022) Chinese 0.61, 0.15 Title, abstract 8 Multi-view learning A 85.21
(Kamateri et al., 2023) CLEFIP-0.54M 0.49, 0.05
Title, abstract, description,
claims, applicants, inventors
731 Ensemble methods A 70.71
Table 4: List of representative papers for patent classification. Note: The first number in dataset size is the sum of training and validation samples, and the second number is the size of the test set. The unit here is million. The metric P stands for precision, while A stands for accuracy. There are multiple results from each paper, and methods with the best performance are selected.
In addition, some research studied variations of network architecture and optimization methods. Shalaby et al. (2018) improved original paragraph vectors (Le and Mikolov, 2014) to represent patent documents. The authors used the inherent structure to derive a hierarchical description of the document, which was more suitable for capturing patent content. In addition, Abdelgawad et al. (2019) analyzed hyper-parameter optimization methods for different neural networks, such as CNN, RNN, and BERT. The results illustrate that optimized networks could sometimes yield 6% accuracy improvement. Similarly, Zhu et al. (2020) used a symmetric hierarchical convolutional neural network and improved the F1 score by approximately 2% on the Chinese short text patent classification task compared to a conventional convolutional neural network.
Alternatively, images can also serve for automated patent classification, which is introduced in Appendix A.1.
Fine-Tuning Transformer-Based Language Model. Fine-tuning tailors the pre-trained model to the patent classification task. Lee et al. fine-tuned the BERT model for the patent classification task and achieved 81.75% precision (Lee and Hsiang, 2020b), which was more than 7% higher than traditional machine learning under the same setting, using word embedding and classifiers (Li et al., 2018). Moreover, Haghighian et al. (2022) compared multiple transformer-based models, including BERT, XLNet, RoBERTa, and ELECTRA, indicating that XLNet performed the best regarding precision, recall, and F1 score. Furthermore, research found that incorporating further training approaches can optimize the fine-tuning process. For example, Christofidellis et al. (2023) integrated domain-adaptive pre-training and used adapters during fine-tuning, which improved the final classification results by about 2% of F1 score. Transformer-based language models have shown better effectiveness than traditional text embedding and become the mainstream method for text-based problems.
Another study addressed this hierarchical classification task as a sequence generation task. Risch et al. (2020b) implemented the transformer-based models to follow the sequence-to-sequence paradigm, where the input was patent texts and the output was the class, such as F02D 41/02. However, the highest accuracy tested on their datasets was only 56.7%, which demonstrated that the sequence-to-sequence paradigm for patent classification was mediocre.
Hybrid Methods. Hybrid methods refer to combining different approaches to make predictions. TechDoc (Jiang et al., 2022a) is a multimodal deep learning architecture, which synthesizes convolutional, recurrent, and graph neural networks through an integrated training process. Convolutional and recurrent neural networks served to process image and text information respectively, while graph neural networks served for the relational information among documents. This multimodal approach was able to reach greater classification accuracy than previous unimodal methods. The advantage of multimodal models is to leverage different modalities for a comprehensive prediction.
Additionally, Zhang et al. (2022) used a multi-view learning method (Zhao et al., 2017) for patent classification. Multi-view learning integrates and learns from multiple distinct feature sets or views of the same data, aiming to improve model performance by leveraging the complementary information available in different views. In a general multi-view learning pipeline, developers specify a model in each view. They aggregate and train all models collaboratively based on multi-view learning algorithms. Zhang et al. (2022) used the two views of patent titles and patent abstracts and tested multi-view learning on a Chinese patent dataset, demonstrating its effectiveness and reliability.
Moreover, a recent study investigated ensemble models that can combine multiple classifiers for classification (Kamateri et al., 2023). While multi-view learning exploits diverse information from different data sources for a more comprehensive understanding, ensemble methods focus on combining predictions from multiple models to improve prediction accuracy and robustness by reducing the models’ variance and bias, which may all use the same data view. The authors experimented with different ensemble methods and finally achieved 70.71% accuracy, which was higher than using only one classifier.
For a synopsis, we list some representative papers in Table 4, including data sources, dataset size, parts used for training, number of classes, methods, and results. The table shows that researchers use different datasets and metrics to test their models, which complicates the comparison of various methods. Hence, we call for standard benchmarks for patent classification to facilitate the development. In addition, LLMs demonstrate more promising results compared to traditional ML models. However, most research is still based on outdated models, such as GPT-2. The application of recent large-sized LLMs to this task would enhance the effectiveness. Additionally, most research has focused on short texts for patent classification, such as abstracts and claims. Nonetheless, titles and abstracts of patents are generic and do not disclose much relevant information. We recommend future research focus more on detailed patent descriptions that contain detailed and specific information about an invention. Furthermore, domain-adaptive methods that adapt standard LLMs to the patent domain are worth investigating to optimize the performance.
5.2 Patent Retrieval
There are three types of retrieval tasks, prior-art search, patent landscaping, and freedom to operate search. Since previous research mainly focused on prior-art search, we introduce this task first in Section 5.2.1 with its corresponding methods in Section 5.2.2, followed by patent landscaping in Section 5.2.3 and freedom to operate search in Section 5.2.4.
5.2.1 Task Definition of Prior-Art Search
Prior-art search refers to, given a target patent document $X$, automatically retrieving $K$ documents that are the most relevant to $X$ from a patent database. [^12] While most research retrieved patent documents as outputs, prior art can also be other types of published (oral or visual) information. This process is crucial for patent examiners to assess the patentability of a new patent application. Prior-art search is not trivial, due to the intricate patent language and the different terms used in various patent descriptions. In principle, documents from distant fields, including both patent literature and other sources, can compromise the novelty of a new application if they disclose the same combination of patent features. However, these different fields often use distinct nomenclature across all word classes, including nouns, verbs, and adjectives. Additionally, many patents and patent applications create their own terms, which can significantly differ from everyday language and even from the terminology used in the technical field. This creation of terms is not always intentional. Attorneys and patent professionals, who may not be deeply familiar with a specific field, might need precise terms for their descriptions and claims and thus decide to invent names spontaneously. These self-defined terms are often highly artificial and may not appear in any other relevant documents. Such terms do not need to be listed in any dictionary.
Figure 7: Illustration of the patent retrieval process
5.2.2 Methodologies for Prior-Art Search
Researchers have intensively invested in patent retrieval tasks and achieved some encouraging accomplishments. We summarize the general retrieving process in Fig. 7. Data types for further pre-processing include text, metadata, and images. Multiple methods can transform patent data into numerical features, such as the statistical method term frequency-inverse document frequency (TF-IDF), which is widely known from text-based search systems, and deep-learning-based word embedding. Relevance ranking algorithms to retrieve the most relevant documents include best-matching BM25, which is a bag-of-word-type method, and cosine similarity derived from the inner vector product.
We focus on text-based patent retrieval in this section and introduce other methods in Appendix A.2. Traditional methods are keyword-based searching and statistical approaches that can rank document relevance, such as the BM25 algorithm (Robertson et al., 2009). Keyword-based methods refer to a search for exact matches in the target corpus according to the input query. A previous survey summarized all keyword-based methods into three categories, namely query expansion, query reduction, and hybrid methods (Shalaby and Zadrozny, 2019). On the other hand, statistical methods exploit a document’s statistics, such as the frequency of specific terms, to calculate a relevance score based on the inputs. For example, the BM25 algorithm calculates the sore using the occurrence of the query terms in each document. These methods are straightforward but the significant semantic and context information has not been explored. Furthermore, each patent document may use its own nomenclature, which may be defined in the description counter-intuitively to daily use.
Researchers proposed two types of improvement to enhance the patent retrieving performance, from keyword-based to full-text-based and from statistical methods to deep-learning methods. Errors occur inherently in keyword-based methods because different keywords can represent the same technical concepts across various disciplines. Hence, Helmers et al. (2019) adopted entire patent texts for similarity comparison and evaluated various feature extraction methods, such as bag-of-words, Word2Vec, and Doc2Vec. The results demonstrated that full-text similarity search could bring better retrieving quality. Moreover, a large body of research studied different deep learning-based embeddings that transfer texts to numerical values for similarity calculation (Sarica et al., 2019; Hain et al., 2022; Hofstätter et al., 2019; Deerwester et al., 1990; Althammer et al., 2021; Trappey et al., 2021; Vowinckel and Hähnke, 2023). Since word embeddings cannot capture contextual information at a higher level, Hofstätter et al. (2019) improved the Word2Vec model by incorporating global context, yielding up to 5% increase in mean average precision (MAP). At the paragraph level, Althammer et al. (2021) evaluated the BERT-PLI (paragraph-level interactions), which is specifically designed for legal case retrieval (Shao et al., 2020), in both patent retrieval and cross-domain retrieval tasks. However, the authors showed that the performance did not surpass the BM25 baseline and indicated that BERT-PLI was not beneficial for patent document retrieval. In another research, Trappey et al. (2021) trained the Doc2Vec model, an embedding that can capture document-level semantic information, based on patent texts for patent recommendation. Patent recommendation aims to retrieve target patent documents from the database. The results showed that Doc2Vec led to more than 10% improvement compared to bag-of-words methods or word embeddings. This research suggests that document-level embeddings are more promising for document retrieval because they can effectively capture context information.
LLMs have demonstrated effectiveness in text retrieval tasks (Ma et al., 2023). Thus, using LLMs for patent prior-art search is a promising research direction. In addition, studies have shown that integrating retrieval into LLMs can improve factual accuracy (Nakano et al., 2021), downstream task performance (Izacard et al., 2023), and in-context learning capabilities (Huang et al., 2023). These retrieval-augmented LLMs are well-established for handling question-answering tasks (Xu et al., 2024). The application of retrieval-augmented generation in the patent domain is another interesting research direction.
5.2.3 Patent Landscaping
Patent landscaping aims to retrieve patent documents related to a particular topic. Landscaping might have a larger overall strategic value for companies and be more related to the machine learning topic. However, patent landscaping is much less investigated compared to prior-art search.
It is straightforward to relate landscaping to prior-art search in two ways. Firstly, we can consider the target topic as a keyword for patent retrieval and use keyword-based methods to retrieve documents from the database. Alternatively, we can find seed patents to represent the topic and retrieve documents that are related to the seed patents as the result (Abood and Feltenberger, 2018).
Furthermore, researchers developed classification models for patent landscaping, which classify whether a patent belongs to a given topic (Choi et al., 2022; Pujari et al., 2022). Choi et al. (2022) concatenated text embedding of abstracts and graph embedding of subject categories for patent representations. Subsequently, the authors added a simple output layer to conduct the necessary binary classification task. To stimulate the research on patent-landscaping-oriented classification, Pujari et al. (2022) released three labeled datasets with qualitative statistics.
5.2.4 Freedom-to-Operate Search
The freedom-to-operate (FTO) search, also known as the clearance search, is a specific type of patent-related research. The aim is to determine if a particular technology or product would be covered by any active intellectual property rights of another party. This search is critical for companies before launching a new product or service in the market and may be requested by investors as part of a due-diligence process.
This task shares similarities with a prior-art search but has an important difference: The search for prior art takes the technology in question and checks if any prior record (not limited to patent documents or active patents) alone (novelty) or in combination (inventiveness/obviousness) contains all the features of the technology. Thus, the search analyzes if the technology is entirely part of the prior art. The documents or other records of the prior art anticipating the technology may also contain more features in their claims. In contrast, the search for freedom-to-operate checks if there is any active patent (or application still in examination), where an independent claim has fewer features (constituting a more general invention) than the technology in question requires. Thus, individual claims of relevant prior-art documents are practically in their entirety included in the technology. Accordingly, while a technology might be patentable due to its novelty and inventiveness, it could still be covered by an earlier patent or pending application if it incorporates the features of the prior art alongside some additional nonobvious features. Therefore, the new invention would be classified as a more specific dependent invention.
Following the task definition, a freedom-to-operate search needs to analyze the claims of potentially relevant patents (applications) in detail and break down the technology in question. Hence, the automated retrieval of targeted patent claims is the core of this task. Few studies have investigated this type of retrieval task. Freunek and Bodmer (2021) trained BERT for the freedom-to-operate search process. They cut patent descriptions into pieces and use BERT to retrieve relevant claims from a constructed dataset. Their report demonstrated that BERT was able to identify relevant claims in small-scale experiments. As this task is important but widely neglected by the community, we introduce it here and suggest it for future research.
5.3 Information Extraction
5.3.1 Task Definition of Information Extraction
The process of extracting specific information from a text corpus is called information extraction. The goal of information extraction is to transform textual data into a more structured format that can be easily processed for various applications, such as data analysis. Hence, researchers usually use information extraction as a support task for patent analysis. Fig. 8 demonstrates how information extraction is applied in the patent domain. Rule-based and deep-learning-based methods are two main streams of information extraction. The extracted information can be entities, relations, or knowledge graphs, which are constructed based on entities and relations. This information can serve for further tasks, such as patent analysis and patent recommendations for companies.
Figure 8: Demonstration of information extraction process in patent domain
5.3.2 Methodologies for Information Extraction
Rule-Based Methods. Traditional extraction methods are rule-based. Researchers manually pre-define a set of rules and extract desired information based on the rules. Chiarello et al. (2019) designed a rule-based system to extract affordances from patents. For example, one of the rules was “The term user followed by can and adverbs, such as readily efficiently, quickly and easily”. The authors used the extracted results to evaluate the quality of engineering design. Another study devised rules according to the syntactic and lexical features of claims to extract facts (Siddharth et al., 2022a). The authors integrated and aggregated these facts to obtain an engineering knowledge graph, which could support inference and reasoning in various engineering tasks.
Well-defined rules can lead to precise extraction so that this process is transparent and interpretable. However, creating and maintaining rules can be time-consuming, difficult, and biased. Moreover, rule-based methods may struggle with the variability and complexity of natural language.
Deep-Learning-Based Methods. Deep learning methods require labeled datasets that include entities or relations for training. There are different network architectures for model training, such as long short-term memory (LSTM) (Chen et al., 2020) and transformers (Son et al., 2022; Puccetti et al., 2023; Giordano et al., 2022). Notably, Son et al. (2022) stated that most patent analysis research focused on claims and abstracts, neglecting description parts that contain essential technical information. The reason may be the notably larger length of the description, which requires appropriate models that can load such text length. Thus, the authors proposed a framework for information extraction through patent descriptions based on the T5 model.
Deep-learning methods can handle a wide variety of language expressions and are easily scalable with more data and computational power. Moreover, deep learning can capture complex patterns and dependencies in language. Nonetheless, deep learning requires large amounts of high-quality annotated data and computation resources for training.
5.3.3 Applications of Information Extraction
Patent Analysis. Patent analysis involves the analysis of patents with respect to multiple aspects, such as evaluating patent novelty or quality, and forecasting technology trends. For example, Chiarello et al. (2019) extracted sentences with a high likelihood of containing affordance from patents to evaluate the quality of engineering design, for example: “The user can easily navigate a set of visual representations of the earlier views.” Puccetti et al. (2023) identified technologies mentioned in patents to anticipate trends for an accurate forecast and effective foresight.
Patent Recommendation. Patent recommendation refers to suggesting relevant patents to users based on their interests, research, or portfolio. Deng and Ma (2021) extracted the semantic information between keywords in the patent domain and constructed weighted graphs of companies and patents. The authors compared the distance based on weighted graphs to generate recommendations. Chen and Deng (2023) extracted connectivity and quality attributes for pairs of patents and companies. Based on knowledge graphs and deep neural networks, the authors designed an interpretable recommendation model that improved the mean average precision of best baselines by 8.6%.
Engineering Design. Engineering design is a creative process that involves defining a problem, conceptualizing ideas, and implementing solutions. The goal is to develop functional, efficient, and innovative solutions to meet specific requirements. Designers can gain insight by analyzing problems and principal solutions extracted from patent documents. Giordano et al. (2022) adopted transformer-based models to extract technical problems, solutions, and advantageous effects from patent documents and achieved an F1 score of 90%. The extracted information helps reveal valuable information hidden in patent documents and generate novel engineering ideas. Similarly, Jiang et al. (2023c) used pre-trained models to identify the motivation, specification, and structure of inventions with the accuracy of 63%, 56%, 44% respectively compared to expert analysis. From design intent to specific solutions, designers can review patents from a systematic perspective to gain better design insights.
The recent LLMs have shown outstanding capabilities in information extraction, such as in complex scientific texts (Dunn et al., 2022) and medical domain (Goel et al., 2023). Therefore, we anticipate the application of LLMs in the patent field to improve the quality of information extraction.
5.4 Novelty and Inventiveness Prediction
5.4.1 Definition of Patent Feature, Novelty, and Inventiveness
Novelty and inventiveness have a clear legal definition in most jurisdictions, [^13] See Articles 54 and 56 of the European Patent Convention (EPC), Sections 2 and 3 of the UK Consolidated Patent Act, or §102 and §103 of the United States patent law codified in Title 35 Code of Laws (USC) as examples. which may strongly deviate from common associations (European Patent Office, 2000; United States Patent and Trademark Office, 2022). An invention is conceived as a combination of features. It is novel if there is no older document or other form of disclosure [^14] They collectively refer to the prior art or state of the art. that alone includes and/or describes all essential features of the invention. Moreover, an invention may be considered novel for two more reasons. (1) The aggregation of prior disclosures does not reveal all its essential features. (2) Experts in the field would not find it obvious to integrate any missing features, for instance, according to their standard practice within the field.
Thus, the assessment of novelty critically relies on the concept of features. The features are the elements the invention needs to comprise to be the invention and are outlined in the patent claims. The independent claims list the essential features. For example, features can be physical elements and objects (typically nouns together with further specifiers), properties (often adjectives), or processing steps in a method. Importantly, features in many jurisdictions cannot be negative, i.e., a missing property. Some offices allow the exclusion of a specific technology from the prior art through negative limitations if the description clearly states the absence of the feature. Later exclusion of features in the claims based on an absence in the description is not possible. The same feature can have very different names in different documents or even be denoted by a neologism well-defined in corresponding invention descriptions. The high variability of terminology between different documents is a major challenge for the examination and also for LLMs. Therefore, novelty prediction is a well-specified task by law and exhibits a high level of mathematical precision atypical of other language-related tasks. In contrast, inventiveness, determining whether the addition of certain features to existing technology is obvious to an expert, can often be ambiguous.
5.4.2 Task Definition of Novelty and Inventiveness Prediction
Novelty and inventiveness prediction is a binary classification task, aiming to determine whether a new patent is novel, given the existing patent database. [^15] While researchers consider this task as a binary classification task, it may also be useful to assess the level of novelty of an invention. Novelty is one of the essential requirements of patent applications and takes a vast of resources and time for human assessment. In addition, the process of reviewing patents is complex and detail-oriented. Even experienced examiners can overlook critical information or fail in judgment. Automated patent novelty evaluation systems can be used as an auxiliary tool for novelty examination. Therefore, this system is critical, because it can not only improve the quality of patents but also enhance the efficiency of patent examination. Since novelty prediction is substantially based on text analysis, the recent LLMs appear well-suited for this complex task.
Figure 9: Methods for patent novelty prediction
5.4.3 Methodologies for Novelty and Inventiveness Prediction
Indicator-Based Methods. Indicator-based approaches rely on pre-defined indicators to measure patent novelty compared to prior art (Verhoeven et al., 2016; Plantec et al., 2021; Sun et al., 2022; Wei et al., 2024; Schmitt and Denter, 2024). Researchers define indicators from various aspects, such as the citations, and assign each indicator a score or weight based on its importance to calculate novelty scores.
For example, Verhoeven et al. (2016) proposed three dimensions to evaluate technological novelty, including novelty in recombination, novelty in technological origins, and novelty in scientific origins. They involved patent classification codes and citation information to analyze each indicator, showing that technological novelty in each dimension was interrelated but conveyed different information. Additionally, Plantec et al. (2021) investigated technological originality, which was defined as the degree of divergence between underlying knowledge couplings embedded in the invention and the predominant design. The authors used proximity indicators of direct citations, co-citation, cosine similarity, co-occurrence, and co-classification, with normalization methods to balance different indicators.
Notably, most research focused on evaluating technological novelty and originality, which was different from the formal definition of patent novelty. While patent novelty is assessed based on prior art and disclosures in a legal context, technological novelty and originality are analyzed in a technical context, focusing on the advancement and uniqueness of the technological contribution. In addition, indicator-based methods also pose some limitations. The selection and weighting of indicators can be subjective, and the indicator may not fully capture the nuanced aspects of inventions.
Outlier Detection. Outlier detection is based on the assumption that novel inventions can be seen as outliers within the landscape of existing patents (Wang and Chen, 2019; Zanella et al., 2021; Jeon et al., 2022). Researchers used text embeddings to represent patents and applied outlier detection algorithms to identify patents that are different from the majority.
[…]